1.Suppose that the manufacturer of a gas clothes dryer has found that when the unit price is p​ dollars, the revenue R​ (in dollars) is R(p)=−5p2+15,000p.

​(a) At what prices p is revenue​ zero?

​(b) For what range of prices will revenue exceed $1,500,000​?

2. A ball is thrown vertically upward with an initial velocity of 80 feet per second. The distance s​ (in feet) of the ball from the ground after t seconds is s=80t-16t2.

​(a) At what time t will the ball strike the​ ground?

​(b) For what time t is the ball more than 84 feet above the​ ground?
Answer #1

4) when the uni price of dryer is f dollars. then revenue (R) in dollars is R.(P) -sp2 +15,ooop (a) Revenue is zero: R(P)=0 =

answered by: ANURANJAN SARSAM
Answer #2

The solutions are given below

Question -1) Given that RCD) = -5.52+ 15000 6 (a) RSb) = -5.62 +15000 $ -5b1b-3000) 56. b-3000 =0 p=t = 3ooo Thuss at price pThit the =0 Question=2) s(t) = got- 16th (a) For ball to s(t) ground Qot-16th 20 16t(5-t) o >tzo and 5-tzo st and ta5 so the